# Goal-guided Generative Prompt Injection Attack on Large Language Models

### v1

@@ 3.3 Solution of Problem (8) and Cos Similarity
where $x$ (known) and $x^{\prime}$ (unknown) are the embedded representations of clean input and adversarial input respectively. Note that in our implementation, we relax the constraint satisfaction problem of the optimal solution $x^{\prime*}$ as
@@
From the perspective of a black-box attack, when $\cos(x^{\prime},x)=1$, then the vectors $x$ and $x^{\prime}$ will have the same direction. Since in NLP, when using vectors (often using unit vectors) to represent text, we care more about the direction of the vector, so $x=x^{\prime}$. In addition, note that $w(\cdot)$ is a bijective function, then for clean text $t$ and adversarial text $t^{\prime}$, there is $t=t^{\prime}$. There we have $r=M(t)=M(t^{\prime})=r^{\prime}$, which contradicts the condition $r\neq r^{\prime}$ in problem (1). When $\cos(x^{\prime},x)=0$, then $w(t^{\prime})$ and $w(t)$ are linearly uncorrelated, which is conflicted with the condition $D(t^{\prime},t)=0$ in problem (1).
3.4 Goal-guided Generative Prompt Injection Attack
@@
Note that $w(t)=x$ and $w(t^{\prime})=x^{\prime}$, based on the previous discussion we can simplify our problem (1) into the following constraint satisfaction problem (CSP)
@@
where $x$ and $x^{\prime}$ represent clean input (known) and adversarial input (unknown) respectively, while $w(\cdot)$ represents the embedded representation of the text (literal meaning) and $\mathcal{D}(\cdot,\cdot)$ represents the distance between the semantics (intrinsic meaning) of the two texts. The hyperparameters $\delta$ and $\epsilon$ are used to control the difficulty of searching constraint, where $\delta$ or $\epsilon$ is smaller, the search accuracy is higher.
@@
Next, we first find the core word set that determines the text semantics through the semantic constraint condition (15), and then use the core word set to generate adversarial text that satisfies the cos similarity condition (15). It is worth noting that the embedded representation is defined on texts, so we use the BERT model to convert any text t into $w(t)$. However, in our work, the semantic distance between texts is defined on the core words of the text, so we use the word2vec model to define the semantic distance $\mathcal{D}(t^{\prime},t)$ between text $t^{\prime}$ and $t$.
3.5 Solving Semantic Constraint(15)
@@
In a text paragraph, usually the first sentence is a summary of the entire paragraph (may be some exceptions that we ignore), where the meaning of a text will be determined by its subject, predicate and object. Therefore, the subject $S_{t}$, predicate $P_{t}$ and object $O_{t}$ that appear first in text $t$ will serve as the core words of the text ($n=3$)
@@
3.6 Solving Cos Similarity Constraint (15)
Next, we will generate adversarial text that satisfies constraint (15) through the core vocabulary $C(t^{\prime})$ of adversarial text. Note that in order to increase the randomness of the sentence, we add another random number $N_{t^{\prime}}$ between $10$ and $100$ as the core word. Now the core vocabulary of the adversarial text $t^{\prime}$ becomes ($n=4$)
$C(t^{\prime})=\{\omega_{1}(t^{\prime})=S_{t^{\prime}},\quad\omega_{2}(t^{\prime})=P_{t^{\prime}},\quad\omega_{3}(t^{\prime})=O_{t^{\prime}},\quad\omega_{4}(t^{\prime})=N_{t^{\prime}}\}.$ (19)
The core word set is embedded into the prompt template to generate a sentence text $t\prime$ with LLM. We iterate $N$ times to randomly generate multiple sentence texts $t^{\prime}$ until the text $t^{\prime}$ satisfies Eqn.(15). Finally, we insert the adversarial text $t^{\prime}$ after the text $t$ to attack the LLM. In fact, it is also feasible to insert $t^{\prime}$ after any sentence in $t$. In Appendix Section A.5, we will see that the difference in attack effectiveness at different locations is minimal.
4 Experiments
4.1 Experimental Details
Below we describe some details of the prompt insertion-based attack method, including the victim model, dataset, and evaluation metrics. In particular, ChatGPT-4-Turbo (gpt-4-0125-preview) is used as our auxiliary model to generate random sentences that comply with grammatical rules. Unless otherwise stated, all results of our algorithm use the parameter settings $\epsilon=0.2$, $\delta=0.05$ and $\gamma=0.5$. We randomly selected 300 examples from the following dataset and tested them using the following two large model families.
4.1.1 Victim Models
  • ChatGPT. ChatGPT is a language model created by OpenAI that can produce conversations that appear to be human-like [22]. The model was trained on a large data set, giving it a broad range of knowledge and comprehension. In our experiments, we choose GPT-3.5-Turbo and GPT-4-Turbo as our victim models in the OpenAI series.
  • Llama-2. Llama-2 [28] is Meta AI’s next-generation open-source large language model. It is more capable than Llama 1 and outperforms other open-source language models on a number of external benchmarks, including reasoning, encoding, proficiency, and knowledge tests. Llama 2-7B, Llama 2-13B and Llama 2-70B are transformer framework-based models.
4.1.2 Q&A Datasets
@@
  • GSM8K The GSM8K dataset is an extensive language model training resource that comprises 800 billion words [1], making it the most massive language dataset available today.
@@
4.1.3 Evaluation Metrics
Assume that the test set is $D$ the set of all question answer pairs predicted correctly by the LLM model $f$ is $T$, and $a(x)$ represents the attack sample generated by the clean input. Then we can define the following three evaluation indicators
@@ 4.1.3 Evaluation Metrics
  • Attack Accuracy The Attack Accuracy metric measures the accuracy on adversarial attack inputs $\mathcal{A}_{\textrm{attack}}=\frac{|\sum_{(x,y)\in D}f(a(x))=y|}{|T|}$.
@@
4.2 Main Results
The experimental results are as shown in Tab. 1. Although we set the insertion position at the end of the clean text, we also give the attack effects of different insertion positions in Appendix Section A.5.
Table 1: Comparison of attack effects of our method G2PIA on seven LLM models and four data sets: including 4 ChatGPT models and 3 Llama models
@@ 4.2 Main Results
The results on four public datasets show that the first-generation ChatGPT-3.5 and ChatGPT-3.5-Turbo have the lowest defence capabilities. Obviously, when ChatGPT first came out, it didn’t think too much about being attacked. Similarly, the small model 7b of Llama-2 is also very weak in resisting attacks. Of course, it is indisputable that the clean accuracy of the models of the Llama series is also very low. The output of small models is more susceptible to noise.
On the other hand, taking ChatGPT-4 as an example, if we compare the ASR values on the 4 data sets, we can conclude that our attack algorithm is more likely to succeed on the data set SQuAD 2.0, while mathematical problems are the most difficult to attack. In contrast to [33] ASR $41.15$ with ChatGPT-3.5 on GSM8k, our attack algorithm with ASR $44.87$ is a general attack strategy and is not specifically designed for problems involving mathematical reasoning.
4.3 Comparison to Other Mainstream Methods
Below we compare our method with the current mainstream black-box attack methods in zero-sample scenarios on two data sets: SQuAD2.0 dataset [23] and Math dataset [11]. Microsoft Prompt Bench [34] uses the following black box attack methods to attack the ChatGPT-3.5 language model, including BertAttack [15], DeepWordBug [5], TextFooler [12], TextBugger [14], Prompt Bench [34], Semantic and CheckList [25]. For the sake of fairness, we also use our method to attack ChatGPT 3.5. Tab. 2 gives the comparison of the results of these methods on the three measurements of Clean Acc, Attack Acc and ASR.
Table 2: Our method is compared with other methods on two datasets
@@ 4.3 Comparison to Other Mainstream Methods
Multiple attack strategies attack ChatGPT-3.5 on two data sets, the SQuAD2.0 dataset and the Math dataset, respectively. As can be seen from Tab. 2, our attack strategy achieves the best results on both data sets. It is worth noting that we count Clean Acc and Attack Acc for each algorithm at the same time, so there are subtle differences between the multiple Clean Acc shown in Tab. 2, but since Clean Acc and Attack Acc are calculated in the same attack algorithm, therefore it has little effect on the value of ASR. Especially on the Math dataset, our algorithm is significantly better than other algorithms, with an ASR of 44.87% compared to BertAttack’s 33.46%. However, our algorithm is a general attack method and is not specifically designed for mathematical problems. To some extent, it is shown that our algorithm has good transfer ability on different types of data sets.
5 Ablation Study
In this section, we will analyze our baseline approach by conducting ablation studies based on two strategies. The first strategy involves extracting sentence components, while the second strategy involves traversing insertion positions. To extract sentence components, we randomly replaced all three components with synonyms. We also randomly performed an ablation study with random breakpoint insertion. The results show that the average ASRs of random location prompt injection and random sentence component replacement multiple times are lower than our method.
Table 3: Ablation studies on datasets of GSM8K and Web-based QA with gpt-3.5-turbo (gpt-3.5-turbo-0125).
@@ 5 Ablation Study
5.1 Transferability
We use the adversarial examples generated by model A to attack model B, thereby obtaining the transferability [33] of the attack on model A. The attack success rate obtained on model B can, to some extent, demonstrate the transferability of the attack on model A. We list the ASR values of the offensive and defensive pairs of LLM into a correlation matrix, as shown in Fig. 4. It can be found that the ChatGPT-4-Turbo attack model has the strongest transferability, while Llama-2-7b has the weakest defensive ability. Obviously, this is because ChatGPT-4-Turbo is the strongest LLMs model among them, while Llama-2-7b is the weakest model. This can be found in the results in Tab. 1.
Figure 4: Transfer Success Rate(TSR) heatmap. The rows and columns represent the attack model and defence model respectively.
5.2 Parameter sensitivity analysis
In our method, the parameters $\epsilon$ and $\delta$ are two important parameters. The former controls the distance between the adversarial text and the clean text in the semantic space, while the latter will affect the optimality of the approximate optimal solution. We selected a total of 9 values from $\{0.1,0.2\cdots,0.9\}$ for the two parameters to attack ChatGPT-3.5 on the GSM8K data set and calculated their ASR values to obtain the curve as shown in Fig. 6 and 6. As shown in Fig. 6, ASR is simply a decreasing function of the distance threshold $\epsilon$. That is, the farther the distance, the worse the attack effect. This aligns with our intuition: injected text that is too far away from the clean text will be treated as noise by LLM and ignored. Fig. 6 shows that when $\gamma=0.5$, our attack strategy achieves the best attack effect. When the gamma value is greater than $0.5$ or less than $0.5$, the attack effect will be attenuated to some extent. It is worth noting that the value of parameter $\gamma$ may vary depending on the model or data. See Appendix sections A.5 and A.6 for more results and discussions.
Figure 5: ASR metric changes with the parameter $\epsilon$
@@ 6 Conclusion
In our work, we propose a new goal-oriented generative prompt injection attack (G2PIA) method. In order to make the injected text mislead the large model as much as possible, we define a new objective function to maximize, which is the KL divergence value between the two posterior conditional probabilities before injection (clean text) and after injection (attack text). Furthermore, we proved that under the condition that the conditional probability follows the multivariate Gaussian distribution, maximizing the KL divergence value is equivalent to maximizing the Mahalanobis distance between clean text and adversarial text. Then we establish the relationship between the optimal adversarial text and clean text. Based on the above conclusions, we design a simple and effective attack strategy with an assisted model to generate injected text that satisfies certain constraints, which will maximize the KL divergence. Experimental results on multiple public datasets and popular LLMs demonstrate the effectiveness of our method.
Appendix 0.A Appendix
0.A.1 Proof on Theorem 1
Assuming that the output value of LLM is a continuous value, below we discuss the maximization problem of $\textrm{KL}(p(y|x),p(y|z))$ under conditional Gaussian distribution
$\max_{\|z\|=1}$ $\textrm{KL}(p(y|x),p(y|z))$ (20)
s.t. $D(s(x),s(z))<\epsilon.$
Let $y$ be an $n\times 1$ random vector and two multivariate Gaussian distributions of $y$ under two different conditions are
$p(y|x)$ $=$ $\mathcal{N}_{1}(y;x,\Sigma)$
$p(y|z)$ $=$ $\mathcal{N}_{2}(y;z,\Sigma)$
Then
$\textrm{KL}(\mathcal{N}_{1}||\mathcal{N}_{2})$ (21)
$=$ $\int\log\frac{\mathcal{N}_{1}(y;x,\Sigma)}{\mathcal{N}_{2}(y;z,\Sigma)}\mathcal{N}_{1}(y;x,\Sigma)\textrm{d}y$
$=$ $\int[-\frac{1}{2}(y-x)^{T}\Sigma^{-1}(y-x)$
$+\frac{1}{2}(y-z)^{T}\Sigma^{-1}(y-z)]\mathcal{N}_{1}(y;x,\Sigma)\textrm{d}y$
$=$ $-\frac{1}{2}\textrm{tr}(\mathbb{E}_{\mathcal{N}_{1}}(y-x)(y-x)^{T}\Sigma^{-1})$
$+\frac{1}{2}\textrm{tr}(\mathbb{E}_{\mathcal{N}_{1}}(y-z)(y-z)^{T}\Sigma^{-1})$
$=$ $-\frac{1}{2}\textrm{tr}(I_{n})+\frac{1}{2}\textrm{tr}\left((z-x)^{T}\Sigma^{-1}(z-x)\right)$
$+\frac{1}{2}\textrm{tr}(\Sigma^{-1}\Sigma)$
$=$ $\frac{1}{2}(z-x)^{T}\Sigma^{-1}(z-x)$
Therefore, assuming that the predicted value of the large model follows a Gaussian distribution, our optimization problem is equivalent to finding a certain text $z$ that maximizes the Mahalanobis distance to $x$. So we optimize the following problems
$\max_{\|z\|_{2}^{2}=1}(z-x)^{T}\Sigma^{-1}(z-x),$ (22)
which is equivalent to the following problem
$\max_{z\neq 0}\frac{(z-x)^{T}\Sigma^{-1}(z-x)}{\|z\|_{2}^{2}}.$ (23)
Further, we transform the above optimization problem as follows
$\min_{z}\|z\|_{2}^{2},\quad s.t.\quad(z-x)^{T}\Sigma^{-1}(z-x)=1,$ (24)
0.A.2 Solution of Problem (8)
Below we first give the form of the solution in a two-dimensional space to explain the problem (8) more intuitively, and then we will give the solution to the problem in the general case.
[…]
Figure 7: When $z^{*}=(z_{1},z_{2})$ is the optimal solution to problem (8), then the vector $\overrightarrow{zo}$ and the normal vector of the ellipse at point $z^{*}$ have the same direction.
For the convenience of discussion, let $\Sigma$ be a diagonal matrix
$\Sigma=\begin{bmatrix}\sigma_{1}^{2}&\\ &\sigma_{2}^{2}\end{bmatrix}$ (25)
The normal vector at point $z=(z_{1},z_{2})$ on the ellipse is
so we have
Finally, we get
Obviously, we can solve the value of $\lambda$ by substituting (29) into Eqn. (27) and $z^{*}$. Note that when $\lambda=0$, then $\lambda$ is on the ellipse, $\|\lambda\|=0$. When $\lambda>0$, then $\lambda$ is outside the ellipse. Otherwise, it is inside the ellipse.
Below we directly derive the relationship between $z$ and $x$ in general based on the KKT condition. The Lagrangian function of the optimization problem is
$\mathcal{L}(z,\lambda)=z^{T}z+\lambda((z-x)^{T}\Sigma^{-1}(z-x)-1).$ (30)
According to the KKT condition, we have
$\frac{\partial\mathcal{L}}{\partial z}=2z+2\lambda\Sigma^{-1}(z-x)=0$ (31)
So
$z^{*}=(\Sigma+\lambda I)^{-1}\lambda x.$ (32)
Finally, if we replace the condition $(\hat{z}-x)^{T}\Sigma^{-1}(\hat{z}-x)=1$ with $(\hat{z}-x)^{T}\Sigma^{-1}(\hat{z}-x)\leq 1$, then we have Lagrange multiplier $\lambda\geq 0$, that is
0.A.3 Relationship between KL Divergence and Cos Similarity in 2-dimensional space
When point $z$ moves on the ellipse, its length corresponds to the solution of the optimization problem: when it takes the minimum value, the optimization problem (20) takes the maximum value. It can be seen from Fig. 7 that when pointing $z$ moves from $A$ to $B$, $\cos(x^{\prime},x)$ gradually increases, but the length of $z$ first decreases and then increases. This process corresponds to that as the cos similarity increases, the KL divergence value first increases from a non-optimal value to the maximum value, and then decreases. We experimentally observed the attack effect as shown in Fig. 8, which also verified our conclusion.
Figure 8: The measurement ASR varies with the changing parameter $\gamma$.
0.A.4 Discussion on KL Divergence and Cos Similarity
In this section, we will explain the reasons for using cos similarity instead of KL divergence from another perspective.
We expand the KL divergence distance function with the first-order Taylor expansion at $x$ to get
$f(x^{\prime})=\textrm{KL}(p(y|x),p(y|x,x^{\prime}))$ (34)
$\approx$ $\textrm{KL}(p(y|x),p(y|x))+\nabla_{x^{\prime}}f(x^{\prime})^{T}(x^{\prime}-x)$
$=$ $\|\nabla_{x^{\prime}}f(x^{\prime})\|_{2}\cos\theta\|x^{\prime}-x\|_{2},$
$\propto$ $\|x^{\prime}-x\|_{2}$
where $\theta$ is the angle between the gradient $\nabla_{x^{\prime}}f(x^{\prime})$ and $x^{\prime}-x$.
In our problem, we assume that $z=\begin{bmatrix}x&x^{\prime}\end{bmatrix}^{T}$ has unit length, so we have
$\|x\|_{2}^{2}+\|x^{\prime}\|_{2}^{2}=1.$ (35)
There are further
$\|x^{\prime}-x\|_{2}^{2}$ $=$ $\|x^{\prime}\|_{2}^{2}+\|x\|_{2}^{2}-2\|x^{\prime}\|_{2}\|x\|_{2}\cos(x^{\prime},x)$ (36)
$=$ $1-2\|x^{\prime}\|_{2}\|x\|_{2}\cos(x^{\prime},x)$
Therefore, we can choose $\cos(x^{\prime},x)$ as the surrogate objective function of KL divergence.
0.A.5 Discussion on Insertion Position of Injection Text
Although our algorithm is set to always insert the injection text at the end of the normal text, from our experiments, the insertion position of the injection text has minimal impact on the attack performance. Tab. 4 takes the most difficult-to-attack problem, such as the Math dataset, as an example and lists the prediction results of inserting injection text into the same normal text at different positions, where the injection text is highlighted in green. In this question, the output of this normal text should be the number $105$. It can be seen that although the position of the injection text affects the output of LLM, these outputs are all wrong answers. This also shows that our algorithm generates injection text that has a strong ability to interfere with normal text, and the effect of the attack is not affected by the position of the injection text.
Table 4: Impact of the insertion position of the injected text on the attack effect: Different insertion positions of LLM output different results, but they are not equal to the correct answer.

### v2

@@ III-C Solving Problem (7) Approximately via Cos Similarity
where $x$ (known) and $x^{\prime}$ (unknown) are the embedded representations of clean and adversarial input, respectively. Note that in our implementation, we relax the constraint satisfaction problem of the optimal solution $x^{\prime*}$ as
@@
From the perspective of a black-box attack, when $\cos(x^{\prime},x)=1$, the vectors $x$ and $x^{\prime}$ will have the same direction. When using vectors (often using unit vectors) to represent text, we care more about the direction of the vector, so $x=x^{\prime}$. In addition, note that $w(\cdot)$ is a bijective function, then for clean text $t$ and adversarial text $t^{\prime}$, there is $t=t^{\prime}$. There we have $r=M(t)=M(t^{\prime})=r^{\prime}$, which contradicts the condition $r\neq r^{\prime}$ in problem (1). When $\cos(x^{\prime},x)=0$, then $w(t^{\prime})$ and $w(t)$ are linearly uncorrelated, which is conflicted with the condition $D(t^{\prime},t)=0$ in problem (1).
III-D Goal-guided Generative Prompt Injection Attack
@@
Note that $w(t)=x$ and $w(t^{\prime})=x^{\prime}$. Based on the previous discussion, we can simplify our problem (1) into the following constraint satisfaction problem (CSP)
@@
where $x$ and $x^{\prime}$ represent clean input (known) and adversarial input (unknown), respectively, while $w(\cdot)$ represents the embedded representation of the text (literal meaning) and $\mathcal{D}(\cdot,\cdot)$ represents the distance between the semantics (intrinsic meaning) of the two texts. The hyperparameters $\delta$ and $\epsilon$ are used to control the difficulty of searching constraint, where $\delta$ or $\epsilon$ is smaller, the search accuracy is higher.
@@
Next, we first find the core word set that determines the text semantics through the semantic constraint condition (14) and then use the core word set to generate adversarial text that satisfies the cos similarity condition (14). It is worth noting that the embedded representation is defined on texts , so we use the BERT model to convert any text $t$ into $w(t)$. However, in our work, the semantic distance between texts is defined on the core words of the text, so we use the word2vec model to define the semantic distance $\mathcal{D}(t^{\prime},t)$ between text $t^{\prime}$ and $t$.
III-E Solving Semantic Constraint(14)
@@
In a text paragraph, usually, the first sentence is a summary of the entire paragraph (maybe some exceptions that we ignore), where the meaning of a text will be determined by its subject, predicate, and object. Therefore, the subject $S_{t}$, predicate $P_{t}$ and object $O_{t}$ that appear first in text $t$ will serve as the core words of the text ($n=3$)
@@
III-F Solving Cos Similarity Constraint (14)
Next, we will generate adversarial text that satisfies constraint (14) through the core vocabulary $C(t^{\prime})$ of adversarial text. Note that to increase the randomness of the sentence, we add another random number $N_{t^{\prime}}$ between $10$ and $100$ as the core word. Now the core vocabulary of the adversarial text $t^{\prime}$ becomes ($n=4$)
$\begin{split}C(t^{\prime})=\{&\omega_{1}(t^{\prime})=S_{t^{\prime}},\omega_{2}(t^{\prime})=P_{t^{\prime}},\\ &\omega_{3}(t^{\prime})=O_{t^{\prime}},\omega_{4}(t^{\prime})=N_{t^{\prime}}\}.\end{split}$ (18)
The core word set is embedded into the prompt template to generate a sentence text $t^{\prime}$ with LLM. We iterate $N$ times to randomly generate multiple sentence texts $t^{\prime}$ until the text $t^{\prime}$ satisfies Eqn. (14). Finally, we insert the adversarial text $t^{\prime}$ after the text $t$ to attack the LLM. Inserting $t^{\prime}$ after any sentence in $t$ is also feasible. In Appendix Section G, we will see minimal difference in attack effectiveness at different locations.
IV Experiments
IV-A Experimental Details
Below we describe some details of the prompt insertion-based attack method, including the victim model, dataset, and evaluation metrics. In particular, ChatGPT-4-Turbo (gpt-4-0125-preview) is used as our auxiliary model to generate random sentences that comply with grammatical rules. Unless otherwise stated, all results of our algorithm use the parameter settings $\epsilon=0.2$, $\delta=0.05$ and $\gamma=0.5$. We randomly selected 300 examples from the following dataset and tested them using two large model families.
IV-A1 Victim Models
  • ChatGPT. ChatGPT is a language model created by OpenAI that can produce conversations that appear human-like [25]. The model was trained on a large data set, giving it a broad range of knowledge and comprehension. In our experiments, we choose GPT-3.5-Turbo and GPT-4-Turbo as our victim models in the OpenAI series.
  • Llama-2. Llama-2 [26] is Meta AI’s next-generation open-source large language model. It is more capable than Llama 1 and outperforms other open-source language models on several external benchmarks, including reasoning, encoding, proficiency, and knowledge tests. Llama 2-7B, Llama 2-13B, and Llama 2-70B are transformer framework-based models.
IV-A2 Q&A Datasets
@@
  • GSM8K The GSM8K dataset, consisting of 800 billion words [27], is the largest language model training resource available today.
@@
IV-A3 Evaluation Metrics
Assume that the test set is $D$, the set of all question answer pairs predicted correctly by the LLM model $f$ is $T$, and $a(x)$ represents the attack sample generated by the clean input. Then we can define the following three evaluation indicators
@@ IV-A3 Evaluation Metrics
  • Attack Accuracy The Attack Accuracy metric measures the accuracy of adversarial attack inputs $\mathcal{A}_{\textrm{attack}}=\frac{|\sum_{(x,y)\in D}f(a(x))=y|}{|T|}$.
@@
IV-B Main Results
The experimental results are shown in Tab. I. Although we set the insertion position at the end of the clean text, we also give the attack effects of different insertion positions in Appendix Section G.
TABLE I: Comparison of attack effects of our method G2PIA on seven LLM models and four data sets: including 4 ChatGPT models and 3 Llama models
@@ IV-B Main Results
The results on four public datasets show that the first-generation ChatGPT-3.5 and ChatGPT-3.5-Turbo have the lowest defense capabilities. Obviously, when ChatGPT first came out, it didn’t think too much about being attacked. Similarly, the small model 7b of Llama-2 is also very weak in resisting attacks. Of course, it is indisputable that the clean accuracy of the models of the Llama series is also very low. The output of small models is more susceptible to noise.
On the other hand, taking ChatGPT-4 as an example, if we compare the ASR values on the 4 data sets, we can conclude that our attack algorithm is more likely to succeed on the data set SQuAD 2.0, while mathematical problems are the most difficult to attack. In contrast to ASR $41.15$ with ChatGPT-3.5 on GSM8k in the paper [31], our attack algorithm with ASR $44.87$ is a general attack strategy and is not specifically designed for problems involving mathematical reasoning.
IV-C Comparison to Other Mainstream Methods
Below we compare our method with the current mainstream black-box attack methods in zero-sample scenarios on two data sets: SQuAD2.0 dataset [30] and Math dataset [29]. Microsoft Prompt Bench [32] uses the following black box attack methods to attack the ChatGPT-3.5 language model, including BertAttack [16], DeepWordBug [33], TextFooler [8], TextBugger [6], Prompt Bench [32], Semantic and CheckList [34]. For fairness, we also use our method to attack ChatGPT 3.5. Tab. II compares the results of these methods on the three measurements of Clean Acc, Attack Acc, and ASR.
TABLE II: Our method is compared with other methods on two datasets
@@ IV-C Comparison to Other Mainstream Methods
Multiple attack strategies attack ChatGPT-3.5 on two data sets, the SQuAD2.0 dataset and the Math dataset, respectively. As seen from Tab. II, our attack strategy achieves the best results on both data sets. It is worth noting that we count Clean Acc and Attack Acc for each algorithm at the same time, so there are subtle differences between the multiple Clean Acc shown in Tab. II, but since Clean Acc and Attack Acc are calculated in the same attack algorithm, therefore it has little effect on the value of ASR. Especially on the Math dataset, our algorithm is significantly better than other algorithms, with an ASR of 44.87% compared to BertAttack’s 33.46%. However, our algorithm is a general attack method not specifically designed for mathematical problems. To some extent, it is shown that our algorithm has good transfer ability on different types of data sets.
V Ablation Study
In this section, we will analyze our baseline approach by conducting ablation studies based on two strategies. The first strategy involves extracting sentence components, while the second involves traversing insertion positions. To extract sentence components, we randomly replaced all three components with synonyms. We also randomly performed an ablation study with random breakpoint insertion. The results show that the average ASRs of random location prompt injection and random sentence component replacement multiple times are lower than our method.
TABLE III: Ablation studies on datasets of GSM8K and Web-based QA with gpt-3.5-turbo (gpt-3.5-turbo-0125).
@@ V Ablation Study
V-A Transferability
We use the adversarial examples generated by model A to attack model B, thereby obtaining the transferability [31] of the attack on model A. The attack success rate obtained on model B can, to some extent, demonstrate the transferability of the attack on model A. We list the ASR values of the offensive and defensive pairs of LLM into a correlation matrix, as shown in Fig. 4. It can be found that the ChatGPT-4-Turbo attack model has the strongest transferability, while Llama-2-7b has the weakest defensive ability. ChatGPT-4-Turbo is the strongest LLM model, while Llama-2-7b is the weakest model. This can be found in the results in Tab. I.
Figure 4: Transfer Success Rate(TSR) heatmap. The rows and columns represent the attack model and defense model, respectively.
V-B Parameter sensitivity analysis
In our method, the parameters $\epsilon$ and $\delta$ are two important parameters. The former controls the distance between the adversarial text and the clean text in the semantic space, while the latter will affect the optimality of the approximate optimal solution. We selected a total of 9 values from $\{0.1,0.2\cdots,0.9\}$ for the two parameters to attack ChatGPT-3.5 on the GSM8K data set and calculated their ASR values to obtain the curve as shown in Fig. 6 and 6. As shown in Fig. 6, ASR is simply a decreasing function of the distance threshold $\epsilon$. That is, the farther the distance, the worse the attack effect. This aligns with our intuition: injected text that is too far away from the clean text will be treated as noise by LLM and ignored. Fig. 6 shows that when $\gamma=0.5$, our attack strategy achieves the best attack effect. The attack effect will be somewhat attenuated when the gamma value is greater than $0.5$ or less than $0.5$. It is worth noting that the value of parameter $\gamma$ may vary depending on the model or data. See the Appendix for more results and discussions.
@@ VI Conclusion
In our work, we propose a new goal-oriented generative prompt injection attack (G2PIA) method. To make the injected text mislead the large model as much as possible, we define a new objective function to maximize, which is the KL divergence value between the two posterior conditional probabilities before injection (clean text) and after injection (attack text). Furthermore, we proved that under the condition that the conditional probability follows the multivariate Gaussian distribution, maximizing the KL divergence value is equivalent to maximizing the Mahalanobis distance between clean text and adversarial text. Then, we establish the relationship between the optimal adversarial text and clean text. Based on the above conclusions, we design a simple and effective attack strategy with an assisted model to generate injected text that satisfies certain constraints, maximizing the KL divergence. Experimental results on multiple public datasets and popular LLMs demonstrate the effectiveness of our method.
Appendix A Appendix
A-A Discussion of Conditions or Assumptions of Theorem 1
We cannot confirm whether these assumptions or preconditions hold or not for the LLM model since our algorithm is a black-box attack, and we do not know the form and parameters of the LLM model. However, to ensure that our optimization problem is well-defined, we make reasonable assumptions and constraints on the variables so that these theoretical results can guide our algorithm design:
  • $z$ is a unit vector: If we do not limit the length of $z$ ($\|z\|_{2}$), then our optimization problem will be trivial: our goal is to maximize the Mahalanobis distance between vector $z$ and vector $x$, then $\|z\|_{2}$ tends to infinity, the objective function (Mahalanobis distance) will tend to infinity ($x$ is a constant vector)
    $\lim_{\|z\|_{2}\rightarrow+\infty}(z-x)^{T}\Sigma^{-1}(z-x)$
    $=$ $\lim_{\|z\|_{2}\rightarrow+\infty}\|z\|_{2}^{2}(\frac{z}{\|z\|_{2}}-\frac{x}{\|z\|_{2}})^{T}\Sigma^{-1}(\frac{z}{\|z\|_{2}}-\frac{x}{\|z\|_{2}})$
    $=$ $\lim_{\|z\|_{2}\rightarrow+\infty}\|z\|_{2}^{2}(t^{T}\Sigma^{-1}t)$
    $\geq$ $\lim_{\|z\|_{2}\rightarrow+\infty}\|z\|_{2}^{2}\lambda_{\min}(\Sigma^{-1})=+\infty,$
    where $t=z/\|z\|_{2}$ is a unit vector and $\lambda_{\min}(\Sigma^{-1})$ is the minimum eigenvalue of $\Sigma^{-1}$. Therefore we restrict $z$ to be the unit vector.
  • $z$ is a unit vector: For the embedding vector of text or sentence, we usually care more about the direction of the vector rather than the length, which can be verified from the wide application of cos similarity in natural language processing. Our algorithm design also uses cos similarity distance, essentially the dot product of two unit vectors.
TABLE IV: Impact of the insertion position of the injected text on the attack effect: Different insertion positions of LLM output different results but are not equal to the correct answer.